OpenStudy (anonymous):

The area of a rectangle is 12 square inches. The length is 5 more than twice the width. Find the length and the width.

OpenStudy (mathstudent55):

L = 2W + 5 A = LW = (2W + 5)W = 12 Solve (2W + 5)W = 12 for W Then plug in the value of W into the first equation above to find L.